OpenStudy (anonymous1):

I have a Calculus question concerning the limit of a expression involving logarithm; details in the reply. I have a Calculus question concerning the limit of a expression involving logarithm; details in the reply. @Mathematics

OpenStudy (anonymous1):

I must show that \[\lim_{x\to +\infty} \frac{\ln x}{x} = 0 \] by knowing that \[ \int_{1}^{x} \left( \frac{1}{\sqrt{t}} dt \right) \geq \int_{1}^{x} \left( \frac{1}{t} dt \right) \] Evaluating the integrals above I obtain: \[ 2\sqrt{x} - 2 \geq \ln x \] But I don't know how to continue from here.

OpenStudy (anonymous):

use L hopitals rule since you obtain infinity over infinity from step 1

OpenStudy (anonymous):

what you did just shows a true falsse statement

OpenStudy (anonymous1):

I should do it without using L'Hôpital's rule.

OpenStudy (zarkon):

\[0\le\lim_{x\to +\infty} \frac{\ln x}{x} \le \lim_{x\to +\infty} \frac{2\sqrt{x}-2}{x}=\cdots \]

OpenStudy (anonymous):

0

OpenStudy (anonymous):

is it true?

OpenStudy (anonymous1):

I think I get it now. By Squeeze theorem, right? \[\lim_{x\to +\infty}\frac{2\sqrt{x} - 2}{x}=0\] Then, by Squeeze theorem, \[\lim_{x\to+\infty}\frac{\ln x}{x}=0\]

OpenStudy (zarkon):

yes

OpenStudy (anonymous1):

\[\lim_{x\to+\infty}\frac{2\sqrt{x} - 2}{x} = \lim_{x\to+\infty}\frac{2\frac{1}{\sqrt{x}} - \frac{2}{x}}{1}=0\]